Operations Report 4-Sept-2021

Work Party Crew Operations Report 04-Sept-2021

This is one of two operations reports filed today


Name of person filing report: Shannon Rupert

Non-nominal systems: Intern trailer water, intern trailer heat.

Notes on non-nominal systems: Working on repairs

Generator: Nothing to report

Hours run: N/A

Solar— SOC % Last 24 hours :

Average 85.2

Minimum 61

Maximum 100

Station Propane Reading – 80 %

Ethanol Free Gasoline – not checked gallons

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 200 gallons

Static to Loft Pump used – yes

Water in GreenHab –0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: Yes

Hours: 216.4

Beginning charge: 96

Ending charge: 97

Currently charging: yes

Notes: When we first arrived, Percy was parked outside my trailer and plugged in but the malfunction light was flashing and it was at 18% when it should have been at 100%. We couldn’t determine why it was not charging so we moved it to the RAM and tried to charge it again. It is slowly charging and the light has stopped blinking. Later we discovered that the 110 plug in the outpost power tower was tripped, probably due to heavy rain. Rover appears to be nominal now.

Notes on rovers: The 4 remaining rovers have had their batteries replaced and were supposed to be delivered this morning, but there was a lack of communication between parties and they did not get delivered. My assumption is that delivery will be rescheduled for sometime after the holiday weekend.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). The three 350’s. The Honda is still non-operational. The 300 is no longer at MDRS.

Reason for use: Test for battery and road worthiness.

Oil Added? No

ATV Fuel Used: Not checked Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: The 350’s are all functional and in generally good operating order. We had to jump one, and tighten the battery cables on another but they all started. The tires were checked and air added as needed.

HabCar used and why, where? Not until this evening, when the new battery was installed. We will use it tomorrow to get water for the GreenHab,

CrewCar used and why, where? Yes, team took it to the Quarry and it was used to go check out the storage units in town.

Luna used and why, where? Yes, to Grand Junction for work party materials

General notes and comments: Grateful for all three vehicles today

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are as we left them in the station but without batteries, as the batteries were stored over summer in the ScienceDome and we have not put them back in the radios.

Summary of Hab operations: Mostly nominal. We have not turned on the hot water yet, but probably will tonight.

Summary of GreenHab operations: Potting soil was purchased We will add soil to the GreenHab. Fill the water tank and put up the shade cloth tomorrow. Seeds will be planted within the next couple of days.

Summary of ScienceDome operations: All nominal, no work done here.

Summary of RAM operations: This building really needs a good solid cleaning.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: We put out poison this summer to avoid rodent infestation, but most of it remained untouched. There was some evidence of pack rat visitors in the intern trailer and the Hab, but remaining buildings show no sign of rodents.

Questions, concerns and requests to Mission Support: Please add Marufa and Ashley to this eList. They are our Operations Interns for this season and they have been copied on this email.

We have requested additional personal from the URC team to assist with the cleaning and mud removal in the storage units.

Copyright © The Mars Society. All rights reserved. | Main Site